[clas_members] [New Logentry] RC Daily Summary August 16, 2022

yprok at jlab.org yprok at jlab.org
Tue Aug 16 09:55:02 EDT 2022

Logentry Text:
Monday, August 15

Day Shift: Took data with NH3(+), HWP IN. Target polarization ~57%. Bleedthrough tests showed ~1nA current from Hall C.  At noon, 5th pass V  separator was turned on, and the bleedthrough test showed a current of ~0.14nA when Hall C was running at 80 microA. Performed a Moeller run at 4nA -> beam polarization is 83.5%.  After the Moeller run, carbon target was inserted. 

Swing Shift: Took data on C with 4nA.  Special BMT tests were performed simultaneously with data taking. 

Tuesday, August 16th
Owl Shift: Took data on C with 4nA.  
Day Shift. Beam went off at 9:10am for the pass change. Hall B is currently in controlled access. Target group is in the Hall. We plan to replace the carbon target with NH3, take a new TE signal and polarize the sample (+).  Once the beam is back, we will resume production on NH3. 

We will have a RC meeting today at 2 PM.

RC Meeting Coordinates:
Meeting ID: 160 867 4253
Passcode: 786506

Data taking progress:
Day 4.0 nA on NH3 (HWP OUT), completed runs :   16693(69M), 16695(2M)
Swing 4.0 nA on C, completed runs 16697(3.7M), 16698(21M), 16699(10M), 16700(67M), 16701(9.5M)
Owl 4.0 nA on C, completed runs 16702(35.3M), 16704(66M)


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4029753
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220816/2bb72a28/attachment.html>

More information about the clas_members mailing list